Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

PS question a Day

Expert replies

Re: PS 007

by wlvoh » Mon Feb 19, 2007 12:58 pm
axefx wrote:
rajs.kumar wrote:PS 007

What is the probability that a card selected from a deck will be either an ace or a spade?
A. 2/52
B. 2/13
C. 7/26
D. 4/13
E. 17/52
Answer is E. 17/52
4 aces and 13 spade cards... this is an OR condition so add the two probabilities:

4/52 + 13/52 = 17/52 !

I'm not so sure your logic here is correct. Imagine a deck of just four cards, an Ace and King of spades, and an Ace and King of clubs. Now, given the same question, "what is the probability of drawing either an ace or a spade?"

By your logic:

Probability of drawing an Ace = 2 / 4 = 1/2.

Probability of drawing a spade = 2 / 4 = 1/2.

Added together, 1/2 + 1/2 is 1. In other words, that says the probability that you will draw either an Ace or a spade is 100%. That clearly CANNOT be the case because I could draw the King of clubs, neither an Ace nor a spade.


I agree with the previous post. You must account for double counting. You have 3 possible cards which satisfy your condition: the two Aces, and the King of spades. Thus, the probability is 3/4.


wlvoh
Join the discussion

by Stacey Koprince » Tue Feb 20, 2007 12:03 am
axefx's first two steps are fine, but you do also have to subtract the overlap.
Please note: I do not use the Private Messaging system! I will not see any PMs that you send to me!!

Stacey Koprince
GMAT Instructor
Director of Online Community
Manhattan GMAT

Contributor to Beat The GMAT!

Learn more about me
Join the discussion

Re: PS004

by Vasudha » Wed Feb 21, 2007 4:20 pm
[the correct answer is 2. It cannot be 3 because 3 and 9 give u 12 and the third term becomes equal to the median 9. therefore when u have 2, the
order is clear viz 2 9 10 and it adds up to 21
Join the discussion

by siddhans » Sun Nov 06, 2011 11:57 am
ajith wrote:PS001


The cost of picture frame M is $10.00 less than 3 times the cost of picture frame N. If the cost of frame M is $50.00, what is the cost of frame N?

A) $13.33
B) $16.66
C) $20.00
D) $26.66
E) $40.00


Answers in three days of time
How do you know if we should write

M = 3N-10 or M = 10 - 3M ??
Join the discussion